\begin{thebibliography}{99}                                                                                               %


\bibitem {carlitz}L. Carlitz, \textit{Degenerate Stirling Bernoulli and
Eulerian numbers,} Utilitas Math. \textbf{15} (1979), 51--88;

\bibitem {Comtet}L. Comtet, \textit{Advanced Combinatorics:} The Art of Finite
and Infinite Expansions, Revised and Enlarged Edition. D. Reidel Publishing
Co., Dordrecht 1974.

\bibitem {dagli}M. C. Da\u{g}l\i, \textit{A new recursive formula arising from
a determinantal expression for weighted Delannoy numbers,} Turkish J. Math
\textbf{45} (2021), no. 1, 471--478; available online at https://doi.org/10.3906/mat-2009-92

\bibitem {dagli1}M. C. Da\u{g}l\i, \textit{Closed formulas and determinantal
expressions for higher-order Bernoulli and Euler polynomials in terms of
Stirling numbers,} Rev. R. Acad. Cienc. Exactas F\'{\i}s. Nat. Ser. A Mat.
RACSAM \textbf{115} (2021), no. 1, Art. No. 32, 8 pages; available online at https://doi.org/10.1007/s13398-020-00970-9

\bibitem {d}L. Dai and H. Pan, \textit{Closed forms for degenerate Bernoulli
polynomials,} Bull. Aust. Math. Soc. \textbf{101} (2020), no. 2, 207--217.

\bibitem {dss}R. Dere, Y. Simsek, and H. M. Srivastava, \textit{A unified
presentation of three families of generalized Apostol type polynomials based
upon the theory of the umbral calculus and the umbral algebra,} J. Number
Theory \textbf{133} (2013), 3245--3263; available online at\ https://doi.org/10.1016/j.jnt.2013.03.004

\bibitem {guo1}B.-N. Guo, D. Lim, and F. Qi, \textit{Maclaurin series
expansions for powers of inverse (hyperbolic) sine, for powers of inverse
(hyperbolic) tangent, and for incomplete gamma functions, with applications to
second kind Bell polynomials and generalized logsine function,} arXiv (2021),
available online at https://arxiv.org/abs/2101.10686v5

\bibitem {guo}B.-N. Guo and F. Qi, \textit{An explicit formula for Bernoulli
numbers in terms of Stirling numbers of the second kind,} J. Anal. Number
Theory \textbf{3} (2015), no. 1, 27--30.

\bibitem {has}Y. He, S. Araci, and H. M. Srivastava, \textit{Some new formulas
for the products of the Apostol type polynomials, }Adv. Differ. Equ. 2016,
1--18 (2016) (Article ID 287); available online at\ https://doi.org/10.1186/s13662-016-1014-0

\bibitem {hasa}Y. He, S. Araci, H. M. Srivastava, and M. Acikgoz, \textit{Some
new identities for the Apostol--Bernoulli polynomials and the
Apostol--Genocchi polynomials}, Appl. Math. Comput. 262 (2015), 31--41;
available online at\ https://doi.org/10.1016/j.amc.2015.03.132

\bibitem {hu-kim}S. Hu and M.-S. Kim, \textit{Two closed forms for the
Apostol--Bernoulli polynomials,} Ramanujan J. \textbf{46} (2018), no. 1,
103--117; available online at https://doi.org/10.1007/s11139-017-9907-4

\bibitem {kargin}L. Karg\i n, \textit{Some formulae for products of geometric
polynomials with applications,} J. Integer Seq. \textbf{20} (2017), Article
17.4.4, 15 pp.

\bibitem {khan}S. Khan, T. Nahid, and M. Riyasat, \textit{On degenerate
Apostol-type polynomials and applications,} Bolet\'{\i}n de la Sociedad
Matem\'{a}tica Mexicana, \textbf{25} (2019), 509--528; available online at https://doi.org/10.1007/s40590-018-0220-z

\bibitem {ks}N. K\i lar and Y. Simsek, \textit{A new family of Fubini type
numbers and polynomials associated with Apostol--Bernoulli numbers and
polynomials,} J. Korean Math. Soc. \textbf{54} (2017), 1605--1621; available
online at\ https://doi.org/10.4134/JKMS.j160597

\bibitem {ks1}N. K\i lar and Y. Simsek, \textit{Identities and relations for
Fubini type numbers and polynomials via generating functions and }%
$p-$\textit{adic integral approach,} Publ. Inst. Math. (Belgr.) \textbf{106}%
(120) (2019), 113--123; available online at\ https://doi.org/10.2298/PIM1920113K

\bibitem {kim}D. S. Kim, T. Kim, H.-I. Kwon, and J.-W. Park, \textit{Two
variable higher-order Fubini polynomials,} J. Korean Math. Soc. \textbf{55}
(2018), no. 4, 975--986; available online at https://doi.org/10.4134/JKMS.j170573

\bibitem {kim1}T. Kim, D. S. Kim, and G.-W. Jang, \textit{A note on degenerate
Fubini polynomials,} Proc. Jangjeon Math. Soc. \textbf{20} (2017), 521--531;
available online at\ http://dx.doi.org/10.17777/pjms2017.20.4.521

\bibitem {lu}D.-Q. Lu and H. M. Srivastava, \textit{Some series identities
involving the generalized Apostol type and related polynomials,} Comput. Math.
Appl. \textbf{62} (2011), 3591--3602; available online at\ https://doi.org/10.1016/j.camwa.2011.09.010

\bibitem {luo}Q.-M. Luo, \textit{On the Apostol Bernoulli polynomials,}
Central European J. Math. \textbf{2 (}2004), 509--515; available online at https://doi.org/10.2478/BF02475959

\bibitem {luo1}Q.-M. Luo, \textit{Apostol--Euler polynomials of higher order
and Gaussian hypergeometric functions,} Taiwanese J. Math. \textbf{10} (2006),
917--925; available online at\ https://doi.org/10.11650/twjm/1500403883

\bibitem {ls}Q.-M. Luo and H. M. Srivastava, \textit{Some generalizations of
the Apostol--Bernoulli and Apostol--Euler polynomials,} J. Math. Anal. Appl.
\textbf{308 }(2005), 290--302; available online at\ https://doi.org/10.1016/j.jmaa.2005.01.020

\bibitem {ls1}Q.-M. Luo and H. M. Srivastava, \textit{Some relationships
between the Apostol Bernoulli and Apostol Euler polynomials,} Comput. Math.
Appl. \textbf{51} (2006), 631--642; available online at\ https://doi.org/10.1016/j.camwa.2005.04.018

\bibitem {n}N. E. N\"{o}rlund, \textit{Vorlesungen \"{u}ber
Differenzenrechnung,} Springer-Verlag, Berlin, 1924.

\bibitem {qi}F. Qi, \textit{Determinantal expressions and recurrence relations
for Fubini and Eulerian polynomials, }J. Interdiscip. Math. \textbf{22}
(2019), 317--335; Available online at https://doi.org/10.1080/09720502.2019.1624063

\bibitem {qi1}F. Qi, \textit{A determinantal expression and a recursive
relation of the Delannoy numbers,} Acta Univ. Sapientiae Math. \textbf{13}
(2021), no. 1, in press; arXiv prprint (2020), available online at https://arxiv.org/abs/2003.12572

\bibitem {qi3}F. Qi, \textit{An explicit formula for the Bell numbers in terms
of the Lah and Stirling numbers,} Mediterr. J. Math. \textbf{13} (2016), no.
5, 2795--2800; available online at https://doi.org/10.1007/s00009-015-0655-7

\bibitem {qi9}F. Qi, \textit{Derivatives of tangent function and tangent
numbers,} Appl. Math. Comput. \textbf{268} (2015), 844--858; available online
at http://dx.doi.org/10.1016/j.amc.2015.06.123

\bibitem {qi4}F. Qi, V. Cernanova, X.-T. Shi, and B.-N. Guo, \textit{Some
properties of central Delannoy numbers,} J. Comput. Appl. Math. \textbf{328}
(2018), 101--115; available online at https://doi.org/10.1016/j.cam.2017.07.013

\bibitem {qi2}F. Qi and R. J. Chapman, \textit{Two closed forms for the
Bernoulli polynomials,} J. Number Theory, \textbf{159} (2016), 89--100;
Available online at https://doi.org/10.1016/j.jnt.2015.07.021

\bibitem {qi5}F. Qi and B.-N. Guo, \textit{Explicit formulas for special
values of the Bell polynomials of the second kind and for the Euler numbers
and polynomials,} Mediterr. J. Math. 14 (2017), no. 3, Article 140, 14 pages;
available online at https://doi.org/10.1007/s00009-017-0939-1

\bibitem {qi6}F. Qi and B.-N. Guo, \textit{Some Determinantal expressions and
recurrence relations of the Bernoulli polynomials,} Mathematics, \textbf{4}
(2016), no. 4, 1--11.

\bibitem {qi80}F. Qi and B.-N. Guo, \textit{Viewing some ordinary di erential
equations from the angle of derivative polynomials,} Iran. J. Math. Sci.
Inform. 16 (2021), no. 1, in press.

\bibitem {qi15}F. Qi, D. Lim, and B.-N. Guo, \textit{Explicit formulas and
identities for the Bell polynomials and a sequence of polynomials applied to
differential equations,} Rev. R. Acad. Cienc. Exactas F\'{\i}s. Nat. Ser. A
Mat. RACSAM, \textbf{113} (2019), 1--9; available online at https://doi.org/10.1007/s13398-017-0427-2

\bibitem {qi7}F. Qi, D.-W. Niu, and B.-N. Guo, \textit{Some identities for a
sequence of unnamed polynomials connected with the Bell polynomials,} Rev. R.
Acad. Cienc. Exactas F\'{\i}s. Nat. Ser. A Math. RACSAM \textbf{113} (2019),
no. 2, 557--567; available online at https://doi.org/10.1007/s13398-018-0494-z

\bibitem {qi10}F. Qi, D.-W. Niu, D. Lim, and B.-N. Guo, \textit{Closed
formulas and identities for the Bell polynomials and falling factorials,}
Contrib. Discrete Math. \textbf{15}(1) (2020), 163--174; available online at https://doi.org/10.11575/cdm.v15i1.68111

\bibitem {qi12}F. Qi and M.-M. Zheng, \textit{Explicit expressions for a
family of the Bell polynomials and applications,} Appl. Math. Comput.
\textbf{258} (2015), 597--607; available online at https://doi.org/10.1016/j.amc.2015.02.027

\bibitem {skr}S. K. Sharma, W. A. Khan, and C. S. Ryoo, \textit{A parametric
kind of the degenerate Fubini numbers and polynomials,} Mathematics
\textbf{8}(3) (2020), Article no: 405; available online at https://doi.org/10.3390/math8030405

\bibitem {sk}H. M. Srivastava and C. K\i z\i late\c{s}, \textit{A parametric
kind of the Fubini-type polynomials,} Rev. R. Acad. Cienc. Exactas F\'{\i}s.
Nat. Ser. A Mat. RACSAM, \textbf{113} (2019), 3253--3267; available online at https://doi.org/10.1007/s13398-019-00687-4

\bibitem {su}D.-D. Su and Y. He, \textit{Some Identities for the two variable
Fubini polynomials,} Mathematics \textbf{7}(2) (2019), Article no: 115;
available online at https://doi.org/10.3390/math7020115

\bibitem {wei}C.-F. Wei and F. Qi, \textit{Several closed expressions for the
Euler numbers,} J. Inequal. Appl. 2015, \textbf{219} (2015).
\end{thebibliography}

\documentclass[12pt]{article}%
\usepackage{amssymb,latexsym}
\usepackage{amsfonts}
\usepackage{amsmath}
\usepackage[colorlinks=true, pdfstartview=FitV, linkcolor=blue,
citecolor=blue, urlcolor=blue]{hyperref}
\usepackage{color}
\usepackage{amssymb}
\usepackage{graphicx}
\usepackage[compress]{cite}%
\providecommand{\U}[1]{\protect\rule{.1in}{.1in}}
\newtheorem{theorem}{Theorem}

\newtheorem{remark}[theorem]{Remark}

\newenvironment{proof}[1][Proof]{\noindent\textbf{#1.} }{\ \rule{0.5em}{0.5em}}
\numberwithin{equation}{section}
\numberwithin{theorem}{section}
\allowdisplaybreaks
\begin{document}

\title{Degenerate Fubini-type polynomials associated with degenerate
Apostol-Bernoulli and Apostol-Euler polynomials of order $\alpha$}
\author{Muhammet Cihat DA\u{G}LI\\Department of Mathematics, Akdeniz University,\\07058-Antalya, Turkey\\E-mail: mcihatdagli@akdeniz.edu.tr}
\date{}
\maketitle

\begin{abstract}
In this paper, by introducing the degenerate Fubini-type polynomials, we give
several relations with the help of the Fa\`{a} di Bruno formula and some
properties of Bell polynomials, and generating function methods. Also, we
derive some new explicit formulas and recurrence relations for Fubini-type
polynomials and numbers. Associating the degenerate Fubini-type polynomials
newly defined here with degenerate Apostol-Bernoulli polynomials and
degenerate Apostol-Euler polynomials of order $\alpha$ enables us to present
additional relations for some degenerate special polynomials and numbers.

\textbf{Keywords:} Apostol-Bernoulli polynomials, Apostol-Euler polynomials,
degenerate Bernoulli polynomials, Generalized Fubini polynomial, Stirling
numbers, explicit formula, recurrence relation, generating function.

\textbf{Mathematics Subject Classification 2010:} 11B68, 11B37, 05A15, 05A19,
11B83, 11Y55.

\end{abstract}

\section{Introduction}

The higher-order Bernoulli polynomials $B_{n}^{(\alpha)}(x)$ and higher-order
Euler polynomials $E_{n}^{(\alpha)}(x)$, each of degree $n$ in $x$ and in
$\alpha$, are defined by means of the generating functions \cite{n}
\[
\left(  \dfrac{t}{e^{t}-1}\right)  ^{\alpha}e^{xt}=%
{\displaystyle\sum\limits_{n=0}^{\infty}}
B_{n}^{(\alpha)}(x)\dfrac{t^{n}}{n!}%
\]
and%
\[
\text{ }\left(  \dfrac{2}{e^{t}+1}\right)  ^{\alpha}e^{xt}=%
{\displaystyle\sum\limits_{n=0}^{\infty}}
E_{n}^{(\alpha)}(x)\dfrac{t^{n}}{n!},
\]
respectively. For $\alpha=1$, we have the classical Bernoulli polynomials
$B_{n}(x)$ and Euler polynomials $E_{n}(x),$ defined by means of the following
generating functions:%
\[
\dfrac{te^{xt}}{e^{t}-1}=%
{\displaystyle\sum\limits_{n=0}^{\infty}}
B_{n}(x)\dfrac{t^{n}}{n!}\text{ }\left(  \left\vert t\right\vert <2\pi\right)
\text{ and }\dfrac{2e^{xt}}{e^{t}+1}=%
{\displaystyle\sum\limits_{n=0}^{\infty}}
E_{n}(x)\dfrac{t^{n}}{n!}\text{ }\left(  \left\vert t\right\vert <\pi\right)
.
\]
In particular, the rational numbers $B_{n}=B_{n}(0)$ and integers $E_{n}%
=2^{n}E_{n}(1/2)$ are called classical Bernoulli numbers and Euler numbers, respectively.

The generalized Apostol-Bernoulli polynomials $B_{n}^{(\alpha)}(x;\gamma)$
were defined by Luo and Srivastava by means of the generating function
\cite{luo,ls,ls1}
\[
\left(  \dfrac{t}{\gamma e^{t}-1}\right)  ^{\alpha}e^{xt}=%
{\displaystyle\sum\limits_{n=0}^{\infty}}
B_{n}^{(\alpha)}(x;\gamma)\dfrac{t^{n}}{n!}%
\]%
\[
\left(  \gamma\in%
\mathbb{C}
;\text{ }\left\vert t\right\vert <2\pi\text{ if }\gamma=1\text{; }\left\vert
t\right\vert <\left\vert \log\gamma\right\vert \text{ if }\gamma\neq1\right)
,
\]
and the generalized Apostol-Euler polynomials $E_{n}^{(\alpha)}(x;\gamma)$ by
means of the generating function \cite{luo1}
\[
\text{ }\left(  \dfrac{2}{\gamma e^{t}+1}\right)  ^{\alpha}e^{xt}=%
{\displaystyle\sum\limits_{n=0}^{\infty}}
E_{n}^{(\alpha)}(x;\gamma)\dfrac{t^{n}}{n!}%
\]%
\[
\left(  \left\vert t\right\vert <\pi\text{ if }\gamma=1\text{; }\left\vert
t\right\vert <\left\vert \log(-\gamma)\right\vert \text{ if }\gamma
\neq1;\text{ }1^{\gamma}=1\right)  .
\]
Carlitz \cite{carlitz} defined degenerate Bernoulli polynomials and degenerate
Euler polynomials by%
\[
\frac{t}{\left(  1+\lambda t\right)  ^{1/\lambda}-1}\left(  1+\lambda
t\right)  ^{x/\lambda}=%
{\displaystyle\sum\limits_{n=0}^{\infty}}
B_{n}(x;\lambda)\dfrac{t^{n}}{n!},
\]
and%
\[
\frac{2}{\left(  1+\lambda t\right)  ^{1/\lambda}+1}\left(  1+\lambda
t\right)  ^{x/\lambda}=%
{\displaystyle\sum\limits_{n=0}^{\infty}}
E_{n}(x;\lambda)\dfrac{t^{n}}{n!}.
\]
For $x=0,$ these are called as degenerate Bernoulli and Euler numbers.

The degenerate versions of Apostol-Bernoulli polynomials and Apostol-Euler
polynomials of order $\alpha$ were introduced by \cite{khan}%
\begin{equation}
\left(  \frac{t}{\gamma\left(  1+\lambda t\right)  ^{1/\lambda}-1}\right)
^{\alpha}\left(  1+\lambda t\right)  ^{x/\lambda}=%
{\displaystyle\sum\limits_{n=0}^{\infty}}
B_{n}^{\left(  \alpha\right)  }(x;\lambda;\gamma)\dfrac{t^{n}}{n!} \label{19}%
\end{equation}
and%
\begin{equation}
\left(  \frac{2}{\gamma\left(  1+\lambda t\right)  ^{1/\lambda}+1}\right)
^{\alpha}\left(  1+\lambda t\right)  ^{x/\lambda}=%
{\displaystyle\sum\limits_{n=0}^{\infty}}
E_{n}^{(\alpha)}(x;\lambda;\gamma)\dfrac{t^{n}}{n!}, \label{21}%
\end{equation}
respectively. Note that since $\lim_{\lambda\rightarrow0}\left(  1+\lambda
t\right)  ^{1/\lambda}=e^{t}$, for $\lambda\rightarrow0,$ $\alpha=\gamma=1,$
the equations (\ref{19}) and (\ref{21}) reduce to the generating functions for
classical Bernoulli and Euler polynomials, respectively.

Let us mention that the above polynomials have been discussed detailed in the
literature. (See for example \cite{dss,has,hasa,lu,luo1} and related
references therein).

We now focus on K\i lar and Simsek's recent study \cite{ks}, in which a family
of Fubini-type polynomials $a_{n}^{\left(  \alpha\right)  }\left(  x\right)  $
are introduced as in the following%
\begin{equation}
\frac{2^{\alpha}}{\left(  2-e^{t}\right)  ^{2\alpha}}e^{xt}=%
{\displaystyle\sum\limits_{n=0}^{\infty}}
a_{n}^{\left(  \alpha\right)  }\left(  x\right)  \frac{t^{n}}{n!},\text{
\ \ }\alpha\in%
\mathbb{N}
_{0}\text{ and }\left\vert t\right\vert <\log2. \label{0}%
\end{equation}
In particular, $a_{n}^{\left(  \alpha\right)  }\left(  0\right)
=a_{n}^{\left(  \alpha\right)  }$ are called Fubini-type numbers. They gave
some relationships between these polynomials and numbers, and other celebrated
polynomials and numbers such as Apostol-Bernoulli numbers, the Frobenius-Euler
numbers and the Stirling numbers via generating function methods and
functional equations. Very recently, Srivastava and K\i z\i lates \cite{sk}
extended Fubini-type polynomials $a_{n}^{\left(  \alpha\right)  }\left(
x\right)  $ to parametric kind families of the Fubini-type polynomials by
considering the two special generating functions and obtained many relations
concerning these and other parametric special polynomials and numbers. As
emphasized therein, the Fubini-type polynomials $a_{n}^{\left(  \alpha\right)
}\left(  x\right)  $ are special case of generalized Apostol-Euler polynomials
$E_{n}^{(\alpha)}(x;\gamma).$ More concretely, $E_{n}^{(2\alpha)}%
(x;-1/2)=2^{3\alpha}a_{n}^{\left(  \alpha\right)  }\left(  x\right)  .$

Further investigations on Fubini polynomials and numbers can be found in
\cite{kargin,ks1,kim,qi,skr,su}, and plenty of references cited therein.

On the other hand, Qi and his colleagues have studied a number of explicit and
recursive formulas, and closed forms for some significant polynomials and
numbers by applying the Fa\`{a} di Bruno formula (see Eq. (\ref{2}), below),
some properties of the Bell polynomials of the second kind, and a general
derivative formula for a ratio of two differentiable functions. See
\cite{dagli,dagli1,d,guo,hu-kim,qi1,qi3,qi9,qi4,qi2,qi5,qi6,qi15,qi7,qi12,wei}
and related references.

In this paper, we introduce degenerate version of Fubini-type polynomials as%
\begin{equation}
\frac{2^{\alpha}}{\left(  2-\left(  1+\lambda t\right)  ^{1/\lambda}\right)
^{2\alpha}}\left(  1+\lambda t\right)  ^{x/\lambda}=%
{\displaystyle\sum\limits_{n=0}^{\infty}}
a_{n}^{\left(  \alpha\right)  }\left(  x;\lambda\right)  \frac{t^{n}}%
{n!},\text{ \ \ \ }\lambda\in%
\mathbb{R}
. \label{1}%
\end{equation}
Notice that for $x=0,$ $a_{n}^{\left(  \alpha\right)  }\left(  0;\lambda
\right)  =a_{n}^{\left(  \alpha\right)  }\left(  \lambda\right)  $ are called
degenerate Fubini-type numbers. Also, for $\lambda\rightarrow0$, these reduce
to Fubini-type polynomials $a_{n}^{\left(  \alpha\right)  }\left(  x\right)  $
aforementioned above.

In parallel with the conclusion given in \cite[Remark 4]{sk}, we infer a
relationship between degenerate Fubini-type polynomials and degenerate
Apostol-Euler polynomials of order $\alpha,$ i.e.%
\begin{equation}
a_{n}^{\left(  \alpha\right)  }\left(  x;\lambda\right)  =2^{-3\alpha}%
E_{n}^{(2\alpha)}(x;\lambda;-1/2). \label{22}%
\end{equation}

In this paper, we would like to use Fa\`{a} di Bruno formula and some
properties of Bell polynomials, and generating function methods in order to
obtain some new explicit formulas, closed forms and recurrence relations for
degenerate Fubini-type polynomials and numbers, and Fubini-type polynomials
and numbers. Moreover, we give a relation between degenerate Fubini-type
polynomials and degenerate Apostol-Bernoulli polynomials of order $\alpha$ and
deduce similar formulas for them.

\section{Properties of second kind Bell polynomials}

The Bell polynomials of the second kind $B_{n,k}\left(  x_{1},x_{2}%
,...,x_{n-k+1}\right)  $ for $n\geq k\geq0$ were defined by \cite[p. 134 and
139]{Comtet}
\[
B_{n,k}\left(  x_{1},x_{2},...,x_{n-k+1}\right)  =%
{\displaystyle\sum\limits_{\substack{1\leq i\leq n,\text{ }l_{i}\in\left\{
0\right\}  \cup\mathbb{N}\\{{\textstyle\sum\nolimits_{i=1}^{n}}}%
il_{i}=n,\text{ }{{\textstyle\sum\nolimits_{i=1}^{n}}}l_{i}=k}}^{\infty}}
\frac{n!}{%
{\textstyle\prod\nolimits_{i=1}^{l-k+1}}
l_{i}!}%
{\textstyle\prod\limits_{i=1}^{l-k+1}}
\left(  \frac{x_{i}}{i!}\right)  ^{l_{i}}.
\]
The Fa\`{a} di Bruno formula can be described in terms of the Bell polynomials
of the second kind $B_{n,k}\left(  x_{1},x_{2},...,x_{n-k+1}\right)  $ by%
\begin{equation}
\frac{d^{n}}{dt^{n}}f\circ h\left(  t\right)  =%
{\displaystyle\sum\limits_{k=0}^{n}}
f^{\left(  k\right)  }\left(  h\left(  t\right)  \right)  B_{n,k}\left(
h^{\prime}\left(  t\right)  ,h^{\prime\prime}\left(  t\right)
,...,h^{(n-k+1)}\left(  t\right)  \right)  . \label{2}%
\end{equation}

For $n\geq k\geq0,$ these polynomials satisfy the following relation \cite[p.
135]{Comtet}%
\begin{equation}
B_{n,k}\left(  abx_{1},ab^{2}x_{2},...,ab^{n-k+1}x_{n-k+1}\right)  =a^{k}%
b^{n}B_{n,k}\left(  x_{1},x_{2},...,x_{n-k+1}\right)  , \label{5}%
\end{equation}
where $a$ and $b$ are any complex number. Also, for $n\geq k\geq0,$ the
following formula is valid for the special case of $B_{n,k}$%
\begin{equation}
B_{n,k}\left(  1,1,...,1\right)  =S\left(  n,k\right)  , \label{12}%
\end{equation}
where $S\left(  n,k\right)  $ denotes the Stirling numbers of the second kind,
can be generated by \cite[p. 206]{Comtet}
\[
\frac{\left(  e^{t}-1\right)  ^{k}}{k!}=%
{\displaystyle\sum\limits_{n=k}^{\infty}}
S\left(  n,k\right)  \frac{t^{n}}{n!}.
\]

In \cite[Remark 1]{qi80}, there existed the formula%

\begin{equation}
B_{n,k}\left(  1,1-\lambda,\left(  1-\lambda\right)  \left(  1-2\lambda
\right)  ,...,\prod_{l=0}^{n-k}\left(  1-l\lambda\right)  \right)
=\frac{(-1)^{k}}{k!}%
{\displaystyle\sum\limits_{l=0}^{k}}
\left(  -1\right)  ^{l}\binom{k}{l}\prod_{q=0}^{n-1}\left(  l-q\lambda\right)
, \label{14}%
\end{equation}
which is equivalent to%
\begin{equation}
B_{n,k}\left(  \left\langle \lambda\right\rangle _{1},\left\langle
\lambda\right\rangle _{2},...,\left\langle \lambda\right\rangle _{n-k+1}%
\right)  =\frac{(-1)^{k}}{k!}%
{\displaystyle\sum\limits_{l=0}^{k}}
\left(  -1\right)  ^{l}\binom{k}{l}\left\langle \lambda l\right\rangle _{n},
\label{15}%
\end{equation}
established in \cite[Theorems 2.1 and 4.1]{qi10}. In \cite[Remark 7.5]{guo1},
the explicit formulas (\ref{14}) and (\ref{15}) were rearranged as%
\begin{align*}
&  B_{n,k}\left(  1,1-\lambda,\left(  1-\lambda\right)  \left(  1-2\lambda
\right)  ,...,\prod_{l=0}^{n-k}\left(  1-l\lambda\right)  \right) \\
&  =(-1)^{k}\frac{\lambda^{n-1}(n-1)!}{k!}%
{\displaystyle\sum\limits_{l=1}^{k}}
\left(  -1\right)  ^{l}l\binom{k}{l}\binom{l/\lambda-1}{n-1}%
\end{align*}
for $\lambda\neq0$ and
\begin{equation}
B_{n,k}\left(  \left\langle \lambda\right\rangle _{1},\left\langle
\lambda\right\rangle _{2},...,\left\langle \lambda\right\rangle _{n-k+1}%
\right)  =(-1)^{k}\lambda\frac{(n-1)!}{k!}%
{\displaystyle\sum\limits_{l=1}^{k}}
\left(  -1\right)  ^{l}l\binom{k}{l}\binom{\lambda l-1}{n-1}. \label{17}%
\end{equation}
Here, the generalized binomial coefficient $\binom{z}{w}$ is defined by%
\[
\binom{z}{w}=%
\begin{cases}
\dfrac{\Gamma\left(  z+1\right)  }{\Gamma\left(  w+1\right)  \Gamma\left(
z-w+1\right)  }, & \text{ if }z,w,z-w\in%
\mathbb{C}
-\left\{  -1,-2,...\right\}  \text{;}\\
0, & \text{if }z\in%
\mathbb{C}
-\left\{  -1,-2,...\right\}  \text{ and }w,z-w\in\left\{  -1,-2,...\right\}
\text{.}%
\end{cases}
\]

\section{Results and their proofs}

In this section, we give some computational formulas for degenerate
Fubini-type numbers, some explicit formulas and recurrence relations for
Fubini-type polynomials and numbers and consequently, degenerate
Apostol-Bernoulli polynomials and degenerate Apostol-Euler polynomials of
order $\alpha.$ Also, we present further relations for some polynomials,
considered here.

\begin{theorem}
The degenerate Fubini-type numbers can be computed by the formula:%
\begin{equation}
a_{n}^{\left(  \alpha\right)  }\left(  \lambda\right)  =(n-1)!%
{\displaystyle\sum\limits_{k=1}^{n}}
\frac{\left\langle -2\alpha\right\rangle _{k}}{2^{\alpha+k}}\frac{(-1)^{k}%
}{\lambda^{k-1}k!}%
{\displaystyle\sum\limits_{l=1}^{k}}
(-1)^{l}l\binom{k}{l}\binom{\lambda l-1}{n-1}, \label{23}%
\end{equation}
where $\left\langle x\right\rangle _{n}$ denotes the falling factorial,
defined for $x\in%
\mathbb{R}
$ by
\[
\left\langle x\right\rangle _{n}=\prod\limits_{k=0}^{n-1}\left(  x-k\right)  =%
\begin{cases}
x\left(  x-1\right)  ...(x-n+1), & \text{if }n\geq1\text{;}\\
1, & \text{if }n=0.
\end{cases}
\]
Consequently, for the special case of the degenerate Apostol-Euler polynomials
$E_{n}^{(\alpha)}(x;\lambda;\gamma),$ the following relation holds:%
\begin{equation}
E_{n}^{(2\alpha)}(0;\lambda;-1/2)=(n-1)!%
{\displaystyle\sum\limits_{k=1}^{n}}
\frac{\left\langle -2\alpha\right\rangle _{k}}{2^{k-2\alpha}}\frac{(-1)^{k}%
}{\lambda^{k-1}k!}%
{\displaystyle\sum\limits_{l=1}^{k}}
(-1)^{l}l\binom{k}{l}\binom{\lambda l-1}{n-1}. \label{24}%
\end{equation}

\end{theorem}

\begin{proof}
If we apply $f(u)=(2-u)^{-2\alpha}$ and $u=g(t)=\left(  1+\lambda t\right)
^{1/\lambda}$ to the Fa\`{a} di Bruno formula (\ref{2}), and use (\ref{5}) and
(\ref{17}) then, we find that%
\begin{align}
&  \frac{d^{n}}{dt^{n}}\left(  (2-\left(  1+\lambda t\right)  ^{1/\lambda
})^{-2\alpha}\right) \nonumber\\
&  =%
{\displaystyle\sum\limits_{k=0}^{n}}
\frac{d^{k}}{du^{k}}(2-u)^{-2\alpha}B_{n,k}\left(  \frac{\lambda\left(
1+t\right)  ^{\lambda-1}}{\lambda},\frac{\lambda\left(  \lambda-1\right)
\left(  1+t\right)  ^{\lambda-2}}{\lambda},\right. \nonumber\\
&  ...,\left.  \frac{\lambda\left(  \lambda-1\right)  ...\left(
\lambda-\left(  n-k\right)  \right)  \left(  1+t\right)  ^{\lambda-(n-k+1)}%
}{\lambda}\right)  ,\nonumber\\
&  =%
{\displaystyle\sum\limits_{k=0}^{n}}
\left\langle -2\alpha\right\rangle _{k}(2-u)^{-2\alpha-k}\frac{\left(
1+t\right)  ^{k\lambda-n}}{\lambda^{k}}B_{n,k}\left(  \left\langle
\alpha\right\rangle _{1},\left\langle \alpha\right\rangle _{2}%
,...,\left\langle \alpha\right\rangle _{n-k+1}\right) \nonumber\\
&  =%
{\displaystyle\sum\limits_{k=0}^{n}}
\left\langle -2\alpha\right\rangle _{k}(2-u)^{-2\alpha-k}\frac{\left(
1+t\right)  ^{k\lambda-n}}{\lambda^{k}}(-1)^{k}\lambda\frac{(n-1)!}{k!}%
{\displaystyle\sum\limits_{l=1}^{k}}
\left(  -1\right)  ^{l}l\binom{k}{l}\binom{\lambda l-1}{n-1}. \label{20}%
\end{align}
Now letting $t\rightarrow0$, which is equivalent to $u\rightarrow0$ on both
sides of (\ref{20}) and taking into consideration the generating function for
degenerate Fubini-type numbers (for $x=0$ in equation (\ref{1})) complete the
proof of (\ref{23}). From the relationship (\ref{22}), the identity (\ref{24})
follows readily.
\end{proof}

\begin{theorem}
The Fubini-type polynomials $a_{n}^{\left(  \alpha\right)  }\left(  x\right)
$ possess the explicit formula
\[
a_{n}^{\left(  \alpha\right)  }\left(  x\right)  =2^{\alpha}%
{\displaystyle\sum\limits_{k=0}^{n}}
\binom{n}{k}%
{\displaystyle\sum\limits_{i=0}^{k}}
\left\langle -2\alpha\right\rangle _{i}\left(  -1\right)  ^{i}S(k,i)x^{n-k},
\]
where $S(n,k)$ is the Stirling numbers of the second kind. Also, the
Fubini-type numbers $a_{n}^{\left(  \alpha\right)  }$ can be written in the
form%
\begin{equation}
a_{n}^{\left(  \alpha\right)  }=2^{\alpha}%
{\displaystyle\sum\limits_{i=0}^{n}}
\left\langle -2\alpha\right\rangle _{i}\left(  -1\right)  ^{i}S(n,i).
\label{11}%
\end{equation}
Besides, the generalized Apostol-Euler numbers $E_{n}^{(\alpha)}(x;\gamma)$
can be expressed as
\begin{equation}
E_{n}^{(2\alpha)}(x;-1/2)=2^{4\alpha}%
{\displaystyle\sum\limits_{k=0}^{n}}
\binom{n}{k}%
{\displaystyle\sum\limits_{i=0}^{k}}
\left\langle -2\alpha\right\rangle _{i}\left(  -1\right)  ^{i}S(k,i)x^{n-k}.
\label{25}%
\end{equation}

\end{theorem}

\begin{proof}
From (\ref{2}), (\ref{5}) and (\ref{12}), we have%
\begin{align}
&  \frac{d^{k}}{dt^{k}}\left(  2-e^{t}\right)  ^{-2\alpha}\nonumber\\
&  =%
{\displaystyle\sum\limits_{i=0}^{k}}
\left\langle -2\alpha\right\rangle _{i}\left(  2-e^{t}\right)  ^{-2\alpha
-i}B_{k,i}\left(  -e^{t},-e^{t},...,-e^{t}\right) \nonumber\\
&  =%
{\displaystyle\sum\limits_{i=0}^{k}}
\left\langle -2\alpha\right\rangle _{i}\left(  2-e^{t}\right)  ^{-2\alpha
-i}(-1)^{i}e^{ti}B_{k,i}\left(  1,1,...,1\right) \nonumber\\
&  \rightarrow%
{\displaystyle\sum\limits_{i=0}^{k}}
\left\langle -2\alpha\right\rangle _{i}(-1)^{i}S\left(  k,i\right)  ,\text{
\ \ \ \ as }t\rightarrow0. \label{29}%
\end{align}
Also, it is obvious that $\left(  e^{xt}\right)  ^{(k)}=x^{k}e^{xt}\rightarrow
x^{k},$ as $t\rightarrow0.$ So, by aid of the Leibnitz's formula for the $n$th
derivative of the product of two functions, we get
\begin{align*}
&  \lim_{t\rightarrow0}\frac{d^{n}}{dt^{n}}\left[  \frac{2^{\alpha}}{\left(
2-e^{t}\right)  ^{2\alpha}}e^{xt}\right] \\
&  =2^{\alpha}%
{\displaystyle\sum\limits_{k=0}^{n}}
\binom{n}{k}%
{\displaystyle\sum\limits_{i=0}^{k}}
\left\langle -2\alpha\right\rangle _{i}(-1)^{i}S\left(  k,i\right)  x^{n-k},
\end{align*}
namely, we have $a_{n}^{\left(  \alpha\right)  }\left(  x\right)  $ by
(\ref{0}). For $x=0,$ we immediately arrive at the identity (\ref{11}). The
equation (\ref{25}) can be deduced from the relation between the Fubini-type
polynomials $a_{n}^{\left(  \alpha\right)  }\left(  x\right)  $ and
generalized Apostol-Euler polynomials $E_{n}^{(\alpha)}(x;\gamma),$ given by
(\ref{22}).
\end{proof}

\begin{theorem}
The Fubini-type polynomials $a_{n}^{\left(  \alpha\right)  }\left(  x\right)
$ satisfy the recurrence relation
\[%
{\displaystyle\sum\limits_{k=0}^{n}}
\binom{n}{k}%
{\displaystyle\sum\limits_{i=0}^{n-k}}
\left\langle 2\alpha\right\rangle _{i}(-1)^{i}S\left(  n-k,i\right)
a_{k}^{\left(  \alpha\right)  }\left(  x\right)  =2^{\alpha}x^{n}%
\]
In particular, the Fubini-type numbers $a_{n}^{\left(  \alpha\right)  }$
provide that%
\begin{equation}%
{\displaystyle\sum\limits_{k=0}^{n}}
\binom{n}{k}%
{\displaystyle\sum\limits_{i=0}^{n-k}}
\left\langle 2\alpha\right\rangle _{i}(-1)^{i}S\left(  n-k,i\right)
a_{k}^{\left(  \alpha\right)  }=0. \label{16}%
\end{equation}
In analogy, the generalized Apostol-Euler polynomials $E_{n}^{(\alpha
)}(x;\gamma)$ possess the recurrence relation%
\begin{equation}%
{\displaystyle\sum\limits_{k=0}^{n}}
\binom{n}{k}%
{\displaystyle\sum\limits_{i=0}^{n-k}}
\left\langle 2\alpha\right\rangle _{i}(-1)^{i}S\left(  n-k,i\right)
E_{k}^{(2\alpha)}(x;-1/2)=2^{4\alpha}x^{n}. \label{26}%
\end{equation}

\end{theorem}

\begin{proof}
Since%
\[
\left[  \left(  2-e^{t}\right)  ^{2\alpha}\right]  \left[  \frac{2^{\alpha}%
}{\left(  2-e^{t}\right)  ^{2\alpha}}e^{xt}\right]  =2^{\alpha}e^{xt},
\]
by keeping in mind the generating function of Fubini-type polynomials
(\ref{0}) and by proceeding as in the proof of (\ref{29}), differentiate $n$
times with respect to $t$ on both sides to deduce that%
\begin{align*}
&
{\displaystyle\sum\limits_{k=0}^{n}}
\binom{n}{k}\frac{\partial^{n-k}}{\partial t^{n-k}}\left[  \left(
2-e^{t}\right)  ^{2\alpha}\right]  \frac{\partial^{k}}{\partial t^{k}}\left[
\frac{2^{\alpha}}{\left(  2-e^{t}\right)  ^{2\alpha}}e^{xt}\right] \\
&  =%
{\displaystyle\sum\limits_{k=0}^{n}}
\binom{n}{k}%
{\displaystyle\sum\limits_{i=0}^{n-k}}
\left\langle 2\alpha\right\rangle _{i}\left(  2-e^{t}\right)  ^{2\alpha
-i}(-1)^{i}e^{ti}S\left(  n-k,i\right)  \frac{\partial^{k}}{\partial u^{k}%
}\left[  \frac{2^{\alpha}}{\left(  2-e^{t}\right)  ^{2\alpha}}e^{xt}\right] \\
&  \rightarrow%
{\displaystyle\sum\limits_{k=0}^{n}}
\binom{n}{k}%
{\displaystyle\sum\limits_{i=0}^{n-k}}
\left\langle 2\alpha\right\rangle _{i}(-1)^{i}S\left(  n-k,i\right)
a_{k}^{\left(  \alpha\right)  }\left(  x\right)  ,\text{ \ \ as }%
t\rightarrow0\\
&  =2^{\alpha}x^{n}.
\end{align*}
Setting $x=0$ yields the equation (\ref{16}) immediately. Formula (\ref{26})
can be derived by the same motivation stated in the proofs of our previous theorems.
\end{proof}

\begin{theorem}
The following relationship holds true:%
\begin{equation}
a_{n-2\alpha}^{\left(  \alpha\right)  }\left(  x;\lambda\right)  =\frac
{B_{n}^{\left(  2\alpha\right)  }(x;\lambda;1/2)}{2^{\alpha}\left\langle
n\right\rangle _{2\alpha}}, \label{27}%
\end{equation}
where $B_{n}^{\left(  \alpha\right)  }(x;\lambda;\gamma)$ is the degenerate
Apostol-Bernoulli polynomials of order $\alpha,$ defined by (\ref{19}).
\end{theorem}

\begin{proof}
If we put $\gamma=1/2$ and replace $\alpha$ by $2\alpha$ in (\ref{19}), we
have%
\begin{align*}%
{\displaystyle\sum\limits_{n=0}^{\infty}}
B_{n}^{\left(  2\alpha\right)  }(x;\lambda;1/2)\dfrac{t^{n}}{n!}  &  =\left(
\frac{t}{\dfrac{1}{2}\left(  1+\lambda t\right)  ^{1/\lambda}-1}\right)
^{2\alpha}\left(  1+\lambda t\right)  ^{x/\lambda}\\
&  =2^{\alpha}t^{2\alpha}%
{\displaystyle\sum\limits_{n=0}^{\infty}}
a_{n}^{\left(  \alpha\right)  }\left(  x;\lambda\right)  \frac{t^{n}}{n!}\\
&  =2^{\alpha}%
{\displaystyle\sum\limits_{n=2\alpha}^{\infty}}
a_{n-2\alpha}^{\left(  \alpha\right)  }\left(  x;\lambda\right)  \frac{t^{n}%
}{(n-2\alpha)!}\\
&  =2^{\alpha}%
{\displaystyle\sum\limits_{n=2\alpha}^{\infty}}
\left\langle n\right\rangle _{2\alpha}a_{n-2\alpha}^{\left(  \alpha\right)
}\left(  x;\lambda\right)  \frac{t^{n}}{n!},
\end{align*}
which concludes the proof.
\end{proof}

Let us continue to study degenerate Fubini-type polynomials $a_{n}^{\left(
\alpha\right)  }\left(  x;\lambda\right)  $. Firstly, from the fact%
\begin{align*}
\left(  1+\lambda t\right)  ^{x/\lambda}  &  =%
{\displaystyle\sum\limits_{n=0}^{\infty}}
\left\langle \frac{x}{\lambda}\right\rangle _{n}\lambda^{n}\frac{t^{n}}{n!}\\
&  =%
{\displaystyle\sum\limits_{n=0}^{\infty}}
\left(  x\right)  _{n,\lambda}\frac{t^{n}}{n!},
\end{align*}
where $\left(  x\right)  _{n,\lambda}=x\left(  x-\lambda\right)  ...\left(
x-\left(  n-1\right)  \lambda\right)  $ for $n>0$ with $\left(  x\right)
_{0,\lambda}=1,$ it is easily verify that%
\begin{align*}%
{\displaystyle\sum\limits_{n=0}^{\infty}}
a_{n}^{\left(  \alpha\right)  }\left(  x;\lambda\right)  \frac{t^{n}}{n!}  &
=\frac{2^{\alpha}}{\left(  2-\left(  1+\lambda t\right)  ^{1/\lambda}\right)
^{2\alpha}}\left(  1+\lambda t\right)  ^{x/\lambda}\\
&  =\left(
{\displaystyle\sum\limits_{n=0}^{\infty}}
a_{n}^{\left(  \alpha\right)  }\left(  \lambda\right)  \frac{t^{n}}%
{n!}\right)  \left(
{\displaystyle\sum\limits_{n=0}^{\infty}}
\left(  x\right)  _{n,\lambda}\frac{t^{n}}{n!}\right) \\
&  =%
{\displaystyle\sum\limits_{n=0}^{\infty}}
\left(
{\displaystyle\sum\limits_{k=0}^{n}}
\binom{n}{k}a_{k}^{\left(  \alpha\right)  }\left(  \lambda\right)  \left(
x\right)  _{n-k,\lambda}\right)  \frac{t^{n}}{n!}.
\end{align*}
Comparing the coefficients $\frac{t^{n}}{n!}$ gives the following theorem.

\begin{theorem}
For $n\geq0,$ we have%
\[
a_{n}^{\left(  \alpha\right)  }\left(  x;\lambda\right)  =%
{\displaystyle\sum\limits_{k=0}^{n}}
\binom{n}{k}a_{k}^{\left(  \alpha\right)  }\left(  \lambda\right)  \left(
x\right)  _{n-k,\lambda}.
\]

\end{theorem}

Now, we observe that%
\begin{align*}
&
{\displaystyle\sum\limits_{n=0}^{\infty}}
\left(  a_{n}^{\left(  \alpha\right)  }\left(  x+1;\lambda\right)
-a_{n}^{\left(  \alpha\right)  }\left(  x;\lambda\right)  \right)  \frac
{t^{n}}{n!}\\
&  =\frac{2^{\alpha}\left(  1+\lambda t\right)  ^{x/\lambda}}{\left(
2-\left(  1+\lambda t\right)  ^{1/\lambda}\right)  ^{2\alpha}}\left(  \left(
1+\lambda t\right)  ^{1/\lambda}-1\right) \\
&  =\frac{2^{\alpha}\left(  1+\lambda t\right)  ^{x/\lambda}}{\left(
2-\left(  1+\lambda t\right)  ^{1/\lambda}\right)  ^{2\alpha-1}}\left(
-1+\frac{1}{2-\left(  1+\lambda t\right)  ^{1/\lambda}}\right) \\
&  =\frac{2^{\alpha}\left(  1+\lambda t\right)  ^{x/\lambda}}{\left(
2-\left(  1+\lambda t\right)  ^{1/\lambda}\right)  ^{2\alpha}}-\sqrt{2}%
\frac{2^{\alpha-1/2}\left(  1+\lambda t\right)  ^{x/\lambda}}{\left(
2-\left(  1+\lambda t\right)  ^{1/\lambda}\right)  ^{2\left(  \alpha
-1/2\right)  }}\\
&  =%
{\displaystyle\sum\limits_{n=0}^{\infty}}
a_{n}^{\left(  \alpha\right)  }\left(  x;\lambda\right)  \frac{t^{n}}%
{n!}-\sqrt{2}%
{\displaystyle\sum\limits_{n=0}^{\infty}}
a_{n}^{\left(  \alpha-1/2\right)  }\left(  x;\lambda\right)  \frac{t^{n}}%
{n!}\\
&  =%
{\displaystyle\sum\limits_{n=0}^{\infty}}
\left(  a_{n}^{\left(  \alpha\right)  }\left(  x;\lambda\right)  -\sqrt
{2}a_{n}^{\left(  \alpha-1/2\right)  }\left(  x;\lambda\right)  \right)
\frac{t^{n}}{n!}.
\end{align*}
Comparing the coefficients $\frac{t^{n}}{n!}$ yields the following theorem.

\begin{theorem}
\label{main1}For $n\geq0,$ we have%
\[
a_{n}^{\left(  \alpha\right)  }\left(  x+1;\lambda\right)  =2a_{n}^{\left(
\alpha\right)  }\left(  x;\lambda\right)  -\sqrt{2}a_{n}^{\left(
\alpha-1/2\right)  }\left(  x;\lambda\right)  .
\]

\end{theorem}

Now, if we differentiate both sides of (\ref{1}) with respect to $t,$ we write%
\begin{align}
&  \frac{d}{dt}\left(  \frac{2^{\alpha}\left(  1+\lambda t\right)
^{x/\lambda}}{\left(  2-\left(  1+\lambda t\right)  ^{1/\lambda}\right)
^{2\alpha}}\right) \nonumber\\
&  =2^{\alpha}\left(  \frac{x\left(  1+\lambda t\right)  ^{(x-\lambda
)/\lambda}}{\left(  2-\left(  1+\lambda t\right)  ^{1/\lambda}\right)
^{2\alpha}}+2\alpha\frac{\left(  1+\lambda t\right)  ^{(x-\lambda+1)/\lambda}%
}{\left(  2-\left(  1+\lambda t\right)  ^{1/\lambda}\right)  ^{2\alpha+1}%
}\right)  . \label{3}%
\end{align}
If we replace $x$ by $x_{1}+x_{2}+\lambda$ and evaluate the terms on both
sides of (\ref{3}), separately, then, we have%
\begin{equation}
\frac{d}{dt}\left(  \frac{2^{\alpha}\left(  1+\lambda t\right)  ^{\left(
x_{1}+x_{2}+\lambda\right)  /\lambda}}{\left(  2-\left(  1+\lambda t\right)
^{1/\lambda}\right)  ^{2\alpha}}\right)  =%
{\displaystyle\sum\limits_{n=0}^{\infty}}
a_{n+1}^{\left(  \alpha\right)  }\left(  x_{1}+x_{2}+\lambda;\lambda\right)
\frac{t^{n}}{n!}, \label{6}%
\end{equation}%
\begin{equation}
\left(  x_{1}+x_{2}+\lambda\right)  \frac{2^{\alpha}\left(  1+\lambda
t\right)  ^{(x_{1}+x_{2})/\lambda}}{\left(  2-\left(  1+\lambda t\right)
^{1/\lambda}\right)  ^{2\alpha}}=\left(  x_{1}+x_{2}+\lambda\right)
{\displaystyle\sum\limits_{n=0}^{\infty}}
a_{n}^{\left(  \alpha\right)  }\left(  x_{1}+x_{2};\lambda\right)  \frac
{t^{n}}{n!} \label{7}%
\end{equation}
and%
\begin{equation}
\sqrt{2}\alpha\frac{2^{\alpha+1/2}\left(  1+\lambda t\right)  ^{(x_{1}%
+x_{2}+1)/\lambda}}{\left(  2-\left(  1+\lambda t\right)  ^{1/\lambda}\right)
^{2\left(  \alpha+1/2\right)  }}=\sqrt{2}\alpha%
{\displaystyle\sum\limits_{n=0}^{\infty}}
a_{n}^{\left(  \alpha+1/2\right)  }\left(  x_{1}+x_{2}+1;\lambda\right)
\frac{t^{n}}{n!}. \label{8}%
\end{equation}
Substitute (\ref{6}), (\ref{7}) and (\ref{8}) in (\ref{3}) to reach the
following theorem.

\begin{theorem}
\label{main}For $n\geq0,$ the degenerate Fubini-type polynomials
$a_{n}^{\left(  \alpha\right)  }\left(  x;\lambda\right)  $ satisfy the
recurrence relation%
\begin{align}
&  a_{n+1}^{\left(  \alpha\right)  }\left(  x_{1}+x_{2}+\lambda;\lambda\right)
\nonumber\\
&  =\left(  x_{1}+x_{2}+\lambda\right)  a_{n}^{\left(  \alpha\right)  }\left(
x_{1}+x_{2};\lambda\right)  +\sqrt{2}\alpha a_{n}^{\left(  \alpha+1/2\right)
}\left(  x_{1}+x_{2}+1;\lambda\right)  . \label{18}%
\end{align}
Letting $\lambda\rightarrow0$ and taking $x_{1}+x_{2}=y$ in (\ref{18}) allow
us to derive the following formula for Fubini-type polynomials
\[
a_{n+1}^{\left(  \alpha\right)  }\left(  y\right)  =ya_{n}^{\left(
\alpha\right)  }\left(  y\right)  +\sqrt{2}\alpha a_{n}^{\left(
\alpha+1/2\right)  }\left(  y+1\right)  .
\]

\end{theorem}

\begin{remark}
From the relationships (\ref{27}) and (\ref{22}), the counterpart identities
in Theorems \ref{main1} and \ref{main} can be presented for degenerate
Apostol-Bernoulli polynomials and degenerate Apostol-Euler polynomials of
order $\alpha.$
\end{remark}

\section{Conclusion}

In our recent study, we have introduced and dealt with degenerate version of
Fubini-type polynomials. Utilizing the Fa\`{a} di Bruno formula and some
properties of Bell polynomials, and generating function methods, we have
derived some new explicit formulas, closed forms and recurrence relations for
degenerate Fubini-type polynomials and numbers, and Fubini-type polynomials
and numbers, defined by K\i lar and Simsek \cite{ks}. Furthermore, by
associating the degenerate Fubini-type polynomials with degenerate
Apostol-Bernoulli polynomials and degenerate Apostol-Euler polynomials of
order $\alpha,$ we have presented analog identities for them. As a final note,
a relation involving degenerate Fubini-type polynomials and degenerate
Apostol-Genocchi polynomials of order $\alpha$, defined by \cite[Eq.
2.6]{khan} can be given and further relations can be obtained.

\end{document}